# DiffClass: Diffusion-Based Class Incremental Learning

### v1

@@ Abstract
Class Incremental Learning (CIL) is challenging due to catastrophic forgetting. On top of that, exemplar-free CIL is even more challenging due to forbidden access to data of previous tasks. Recent exemplar-free CIL methods attempt to mitigate catastrophic forgetting by synthesizing previous task data. However, they fail to overcome the catastrophic forgetting due to the inability to deal with the significant domain gap between real and synthetic data. To overcome these issues, we propose a novel exemplar-free CIL method. Our method adopts multi-distribution matching (MDM) diffusion models to align quality of synthetic data and bridge domain gaps among all domains of training data. Moreover, our approach integrates selective synthetic image augmentation (SSIA) to expand the distribution of the training data, thereby improving the model’s plasticity and reinforcing the performance of our multi-domain adaptation (MDA) technique. With the proposed integrations, our method then reformulates exemplar-free CIL into a multi-domain adaptation problem to implicitly address the domain gap problem and enhance model stability during incremental training. Extensive experiments on benchmark CIL datasets and settings demonstrate that our method excels previous exemplar-free CIL methods with non-marginal improvements and achieves state-of-the-art performance.
@@ 1 Introduction
To counteract forgetting knowledge of past tasks, the most recent exemplar-free CIL works [32, 7, 8, 44] propose to synthesize previous data instead of using real data. The synthetic data of previous tasks are generated through either model inversion [42] with knowledge distillation or denoising diffusion models [11]. However, these methods suffer from significant domain gaps between synthetic data and real data especially when the number of incremental tasks is large (i.e. long-term CIL), which inevitably misleads the decision boundaries between new and previous classes. The obtained models favor plasticity over stability, meaning they stubbornly keep previous knowledge in mind without a willingness to learn new knowledge, as demonstrated in Sec. 3. Therefore, how to exhibit both stability and plasticity in exemplar-free CIL remains a crucial challenge.
@@ 2.0.1 Class Incremental Learning (CIL)
has merged as a challenging problem focusing on how models can incrementally learn new classes without forgetting previously acquired knowledge. To overcome the catastrophic forgetting, recent successful approaches [24, 12, 3, 38, 37] store training data from previously learned classes as exemplars and replay them while learning new tasks. Exemplars are indeed helpful for reviewing past task knowledge and thus benefit the incremental learning process. However, due to privacy and legality issues, and memory constraints on devices, it may be unachievable in practice.
@@
In recent developments, diffusion models have also shown remarkable versatility beyond image generation. They are successfully applied in various domains, including general computer vision tasks [39], audio processing [30], and even text-to-3D [22].
@@ 3 Diagnosis: Domain Gaps in Exemplar-Free CIL
Although recent advancements in generative AI can generate realistic images, we notice that the distributions of the generated synthetic images are still different from those of real images with domain gaps, leading to low accuracy in the classes trained with synthetic data in exemplar-free CIL settings. We also further dig into the low accuracy and find that the reason may be the model’s preference for domains over classes after training, i.e. the model classifies whether the image is real or synthetic rather than its true label.
@@
The domain gaps result in unsatisfying performance after model training. In specific, we train a model with only a ResNet [10] backbone and a linear classifier for the first four tasks (each with 5 classes) in a 20-task CIL setting on the ImageNet100 dataset (refer to Sec. 5 for more details). From the second to the fourth tasks, aside from the real data of the current task, we also train with synthetic data of the previous tasks generated by the pre-trained SD V1.5 model. We additionally train another model with entirely real data for the four tasks as a reference for how well the model can perform with real data.
Table 1: Diagnosis experiment accuracy result (in %) of incremental training the model with synthetic previous task data and real data of current task vs. training model with all real data for first two tasks of twenty-task incremental setting on ImageNet100.
@@
In Tab. 1, we present the accuracy on the real test dataset. As observed, the model performs significantly better on the classes of the new task (i.e. class 15-19, trained with real data) than previous tasks (i.e. class 0-14, trained with both real in previous task and synthetic data in the current task), demonstrating the model’s preference for plasticity over stability.
@@ 4.1 Framework
Figure 3: Overview of our model framework learning on currect task $\mathcal{T}_{i+1}$. previous MDM diffusion models $J_{0:i}$ are used to generated Synthetic Data of Previous Tasks $\mathcal{D}_{0:i}^{\text{syn}}$. MDM diffusion model of Current Task is then finetuned using MDM technique using Real Current Task Data $\mathcal{D}_{i}^{\text{real}}$ and randomly sampled small batch of $\mathcal{D}_{0:i}^{\text{syn}}$. $J_{0:i}$ is subsequently used to obtain $\mathcal{D}_{i}^{\text{aug}}$ by SSIA. The model trains with MDA on the combined dataset.
@@ 4.2 Finetuning Multi-Distribution Matching Diffusion Model with LoRA
$\begin{split}\mathcal{L}_{MDM}=&\,||\frac{1}{|\mathcal{D}_{i}^{\text{real}}|+|Z(\mathcal{D}_{0:i}^{\text{syn}})|}\sum_{j=1}^{|\mathcal{D}_{i}^{\text{real}}|+|Z(\mathcal{D}_{0:i}^{\text{syn}})|}\left(\epsilon-\epsilon_{\theta}(x_{t}^{\prime},t)\right)||_{\mathcal{H}}^{2}\\ &\leq\frac{1}{|\mathcal{D}_{i}^{\text{real}}|}\sum_{j=1}^{|\mathcal{D}_{i}^{\text{real}}|}||\epsilon-\epsilon_{\theta}(x_{t},t)||_{\mathcal{H}}^{2}=\mathcal{L}_{diff}\end{split}$ (1)
@@ 4.3 Forming Current Task Training Dataset
    $\mu_{cn}^{\text{gen}}=\frac{1}{|\mathcal{D}_{cn}^{\text{gen}}|}\sum_{\mathbf{x}\in\mathcal{D}_{cn}^{\text{gen}}}\mathbf{x},\text{where }cn\in\mathcal{CN}$ (2)
@@
    $\mathbf{x}_{i}^{cn}\sim\mathcal{N}(\mu_{cn}^{\text{gen}},\text{Cov}_{cn}^{\text{gen}})$ (4)
@@
    $\tau_{cn}^{\text{gen}}=k\cdot\sqrt{\text{diag}(\text{Cov}_{cn}^{\text{gen}})}$ (5)
    $\mathcal{D}_{i}^{\text{aug}}=\bigcup_{cn=1}^{\mathcal{CN}}\{\mathbf{x}_{i}^{cn}\,|\,\|{\mathbf{x}_{i}^{cn}-\mu_{cn}^{\text{gen}}\|\leq\tau_{cn}^{\text{gen}}}\}$ (6)
@@ 4.4 Training with Multi-Domain Adaptation
$\mathcal{L}_{i}^{\text{train}}=\mathcal{L}_{i}^{\text{cls}}+\mathcal{L}_{i}^{\text{da}}$ (7)
@@
$\mathcal{L}_{i}^{\text{cls}}=-\frac{1}{|\mathcal{D}_{i}^{\text{train}}|}\sum_{\mathbf{x}\in\mathcal{D}_{i}^{\text{train}}}y_{c}\log\left(G_{i}\left(F_{i}\left(\mathbf{x}\right)\right)\right)$ (8)
@@
$\mathcal{L}_{i}^{\text{da}}=-\frac{1}{|\mathcal{D}_{i}^{\text{train}}|}\sum_{\mathbf{x}\in\mathcal{D}_{i}^{\text{train}}}\left[y_{d}\log\left(K_{i}\left(F_{i}\left(\mathbf{x}\right)\right)\right)+(1-y_{d})\log\left(1-\left(K_{i}\left(F_{i}\left(\mathbf{x}\right)\right)\right)\right)\right]$ (9)
@@
By reforming exemplar-free CIL as a straightforward task-agnostic multi-domain adaption problem, our method enjoys the following advantages. (i) Our model framework keeps simple without any cumbersome parts, which benefits incremental training efficiency. (ii) More importantly, our model is robust in both stability and plasticity since it is fully capable of learning important feature knowledge from both label classification and domain classification (synthetic vs. real) in each task. (iii) Our proposed method can not only perform well on a test dataset consisting of entirely real data but also elaborate to perform well on entirely synthetic test data and combined image groups (see Appendix , which better simulates the continual learning scenarios in real-world settings.
@@ 5.1 Datasets and Evaluation Protocol
We report both the final accuracy from the last task $Acc_{L}$ and the average incremental accuracy averaged over all incremental tasks ${Acc}_{avg}=\frac{1}{N}\sum_{i=0}^{N-1}Acc_{i}$. Implementation Details. For a fair comparison, for CIFAR100, following previous works [32, 7], we use a modified 32-layer ResNet [10] as the backbone for all approaches. For our model, we train with SGD optimizer for 120 epochs. The learning rate is initially set to 0.1 with a decay factor of 0.1 after 100 epochs. The weight decay is set to 0.0002 and the batch size is 128. For ImageNet100, we use ResNet18 [10] as the backbone for all methods. For our training, the SGD optimizer is adopted to train 40 epochs. The learning rate is initially set to 0.1 with a decay factor of 0.1 after 30 epochs. The weight decay is set to 0.0001 and the batch size is 128. We train all methods from scratch.
@@ 5.2 Results and Analysis
Figure 5: Incremental Accuracy on ImageNet100. Our method greatly outperforms all baseline methods in all incremental settings. Our method achieves more significant improvements in settings with more incremental tasks (such as increasing $N$ from 5 to 10, or from 10 to 20)
@@ 5.3 Ablation Studies
Without finetuning diffusion models with a multi-distribution matching technique, the average accuracy $Acc_{avg}$ drops by 15.14 percent (74.85% vs. 59.71%), and the final classification accuracy $Acc_{N}$ drops by 16.09 percent (67.26% vs. 51.17%). From Fig. 6, we also observe that MDM serves a crucial role in reviewing knowledge of previous tasks (i.e. building model stability).
@@
Without further enhancement from selective synthetic image augmentation, the average accuracy $Acc_{avg}$ drops by 12.48 percent, and the final accuracy $Acc_{L}$ drops by 14.97 percent. Furthermore, Fig. 6 shows that SSIA helps the model not only learn new knowledge (i.e. building model plasticity) but also remember the knowledge from previous tasks.

### v2

@@ Abstract
Class Incremental Learning (CIL) is challenging due to catastrophic forgetting. On top of that, exemplar-free CIL is even more challenging due to forbidden access to data of previous tasks. Recent exemplar-free CIL methods attempt to mitigate catastrophic forgetting by synthesizing previous task data. However, they fail to overcome the catastrophic forgetting due to the inability to deal with the significant domain gap between real and synthetic data. To overcome these issues, we propose a novel exemplar-free CIL method. Our method adopts multi-distribution matching (MDM) diffusion models to align quality of synthetic data and bridge domain gaps among all domains of training data. Moreover, our approach integrates selective synthetic image augmentation (SSIA) to expand the distribution of the training data, thereby improving the model’s plasticity and reinforcing the performance of our multi-domain adaptation (MDA) technique. With the proposed integrations, our method then reformulates exemplar-free CIL into a multi-domain adaptation problem to implicitly address the domain gap problem and enhance model stability during incremental training. Extensive experiments on benchmark CIL datasets and settings demonstrate that our method excels previous exemplar-free CIL methods with non-marginal improvements and achieves state-of-the-art performance. Our project page is available at https://cr8br0ze.github.io/DiffClass.
@@ 1 Introduction
To counteract forgetting knowledge of past tasks, the most recent exemplar-free CIL works [33, 7, 8, 47] propose to synthesize previous data instead of using real data. The synthetic data of previous tasks are generated through either model inversion [45] with knowledge distillation or denoising diffusion models [11]. However, these methods suffer from significant domain gaps between synthetic data and real data especially when the number of incremental tasks is large (i.e. long-term CIL), which inevitably misleads the decision boundaries between new and previous classes. The obtained models favor plasticity over stability, meaning they tend to learn new knowledge but without keeping previous knowledge in mind as demonstrated in Sec. 3. Therefore, how to exhibit both stability and plasticity in exemplar-free CIL remains a crucial challenge.
@@ 2.0.1 Class Incremental Learning (CIL)
has merged as a challenging problem focusing on how models can incrementally learn new classes without forgetting previously acquired knowledge. To overcome the catastrophic forgetting, recent successful approaches [25, 12, 3, 39, 38, 52, 49, 50, 51, 42, 41] store training data from previously learned classes as exemplars and replay them while learning new tasks. Exemplars are indeed helpful for reviewing past task knowledge and thus benefit the incremental learning process. However, due to privacy and legality issues, and memory constraints on devices, it may be unachievable in practice.
@@
In recent developments, diffusion models have also shown remarkable versatility beyond image generation. They are successfully applied in various domains, including other computer vision tasks [40, 19], audio processing [31], and even text-to-3D [23].
@@ 3 Diagnosis: Domain Gaps in Exemplar-Free CIL
Although recent advancements in generative artificial intelligence can generate realistic images, we notice that the distributions of the generated synthetic images are still different from those of real images with domain gaps, leading to low accuracy in the classes trained with synthetic data in exemplar-free CIL settings. We also further dig into the low accuracy and find that the reason may be the model’s preference for domains over classes after training, i.e. the model classifies whether the image is real or synthetic rather than its true label.
@@
These domain gaps can potentially effect model’s performance after model training with real and synthetic data, since the decision boundary can be significantly distorted by synthetic data, as it may treat the real class 0 and class 1 (with a smaller distribution discrepancy) as the same class in testing.
We also conduct an experiment in a class incremental setting to further verify. In specific, we train a model with only a ResNet [10] backbone and a linear classifier for the first four tasks (each with 5 classes) in a 20-task CIL setting on the ImageNet100 dataset (refer to Sec. 5 for more details). From the second to the fourth tasks, aside from the real data of the current task, we also train with synthetic data of the previous tasks generated by the pre-trained SD V1.5 model. We additionally train another model with entirely real data for the four tasks as a reference for how well the model can perform with real data.
Table 1: Diagnosis experiment accuracy result (in %) of incremental training the model with synthetic previous task data and real data of current task vs. training model with all real data for first four tasks of twenty-task incremental setting on ImageNet100.
@@
In Tab. 1, we present the accuracy on the real test dataset at the end of task 4. As observed, the model performs significantly better on the classes of the new task (i.e. class 15-19, trained with real data) than previous tasks (i.e. class 0-14, trained with both real in previous task and synthetic data in the current task), demonstrating the model’s preference for plasticity over stability.
@@ 4.1 Framework
Figure 4: Model Framework Overview learning on currect task $\mathcal{T}_{i+1}$. previous MDM diffusion models $J_{0:i}$ are used to generated Synthetic Data of previous tasks $\mathcal{D}_{0:i}^{\text{syn}}$. MDM diffusion model of current task is then finetuned using MDM technique using Real current task Data $\mathcal{D}_{i}^{\text{real}}$ and randomly sampled small batch of $\mathcal{D}_{0:i}^{\text{syn}}$. $J_{0:i}$ is subsequently used to obtain $\mathcal{D}_{i}^{\text{aug}}$ by SSIA. The model trains with MDA on the combined dataset.
@@ 4.2 Finetuning Multi-Distribution Matching Diffusion Model with LoRA
$\begin{split}\mathcal{L}_{MDM}=&\,||\frac{1}{|\mathcal{D}_{i}^{\text{real}}|+|Z(\mathcal{D}_{0:i}^{\text{syn}})|}\sum_{j=1}^{|\mathcal{D}_{i}^{\text{real}}|+|Z(\mathcal{D}_{0:i}^{\text{syn}})|}\left(\epsilon-\epsilon_{\theta}(x_{t}^{\prime},t)\right)||_{\mathcal{H}}^{2}\\ &\leq\frac{1}{|\mathcal{D}_{i}^{\text{real}}|}\sum_{j=1}^{|\mathcal{D}_{i}^{\text{real}}|}||\epsilon-\epsilon_{\theta}(x_{t},t)||_{\mathcal{H}}^{2}=\mathcal{L}_{diff}.\end{split}$ (1)
@@ 4.3 Forming Current Task Training Dataset
    $\mu_{cn}^{\text{gen}}=\frac{1}{|\mathcal{D}_{cn}^{\text{gen}}|}\sum_{\mathbf{x}\in\mathcal{D}_{cn}^{\text{gen}}}\mathbf{x},\text{where }cn\in\mathcal{CN},$ (2)
@@
    $\mathbf{x}_{i}^{cn}\sim\mathcal{N}(\mu_{cn}^{\text{gen}},\text{Cov}_{cn}^{\text{gen}}),$ (4)
@@
    $\tau_{cn}^{\text{gen}}=k\cdot\sqrt{\text{diag}(\text{Cov}_{cn}^{\text{gen}})},$ (5)
    $\mathcal{D}_{i}^{\text{aug}}=\bigcup_{cn=1}^{\mathcal{CN}}\{\mathbf{x}_{i}^{cn}\,|\,\|{\mathbf{x}_{i}^{cn}-\mu_{cn}^{\text{gen}}\|\leq\tau_{cn}^{\text{gen}}}\}.$ (6)
@@ 4.4 Training with Multi-Domain Adaptation
$\mathcal{L}_{i}^{\text{train}}=\mathcal{L}_{i}^{\text{cls}}+\mathcal{L}_{i}^{\text{da}}.$ (7)
@@
$\mathcal{L}_{i}^{\text{cls}}=-\frac{1}{|\mathcal{D}_{i}^{\text{train}}|}\sum_{\mathbf{x}\in\mathcal{D}_{i}^{\text{train}}}y_{c}\log\left(G_{i}\left(F_{i}\left(\mathbf{x}\right)\right)\right),$ (8)
@@
$\mathcal{L}_{i}^{\text{da}}=-\frac{1}{|\mathcal{D}_{i}^{\text{train}}|}\sum_{\mathbf{x}\in\mathcal{D}_{i}^{\text{train}}}\left[y_{d}\log\left(K_{i}\left(F_{i}\left(\mathbf{x}\right)\right)\right)+(1-y_{d})\log\left(1-\left(K_{i}\left(F_{i}\left(\mathbf{x}\right)\right)\right)\right)\right].$ (9)
@@
By reforming exemplar-free CIL as a straightforward task-agnostic multi-domain adaption problem, our method enjoys the following advantages. (i) Our model framework keeps simple without any cumbersome parts, which benefits incremental training efficiency. (ii) More importantly, our model is robust in both stability and plasticity since it is fully capable of learning important feature knowledge from both label classification and domain classification (synthetic vs. real) in each task. (iii) Our proposed method can not only perform well on a test dataset consisting of entirely real data but also elaborate to perform well on entirely synthetic test data and combined image groups (see Appendix) , which better simulates the continual learning scenarios in real-world settings.
@@ 5.1 Datasets and Evaluation Protocol
We report both the final accuracy from the last task $Acc_{L}$ and the average incremental accuracy averaged over all incremental tasks ${Acc}_{avg}=\frac{1}{N}\sum_{i=0}^{N-1}Acc_{i}$. Implementation Details. For a fair comparison, for CIFAR100, following previous works [33, 7], we use a modified 32-layer ResNet [10] as the backbone for all approaches. For our model, we train with SGD optimizer for 120 epochs. The learning rate is initially set to 0.1 with a decay factor of 0.1 after 100 epochs. The weight decay is set to 0.0002 and batch size of 128. For ImageNet100, we use ResNet18 [10] as the backbone for all methods. For our training, the SGD optimizer is adopted to train 40 epochs. The learning rate is initially set to 0.1 with a decay factor of 0.1 after 30 epochs. The weight decay is set to 0.0001 and batch size of 128. We train and report all methods from scratch with original implementations.
@@ 5.2 Results and Analysis
Figure 8: Incremental Accuracy on ImageNet100. Our method greatly outperforms all baseline methods in all incremental settings. Our method achieves more significant improvements in more incremental task settings (e.g. increase $N$ from 5 to 10 or to 20)
@@ 5.3 Ablation Studies
Without finetuning diffusion models with a multi-distribution matching technique, the average accuracy $Acc_{avg}$ drops by 15.14 percent (74.85% vs. 59.71%), and the final classification accuracy $Acc_{N}$ drops by 16.09 percent (67.26% vs. 51.17%). From Fig. 9, we also observe that MDM serves a crucial role in reviewing previous knowledges (i.e. stability).
@@
Without further enhancement from selective synthetic image augmentation, the average accuracy $Acc_{avg}$ drops by 12.48 percent, and the final accuracy $Acc_{L}$ drops by 14.97 percent. Furthermore, Fig. 9 shows that SSIA helps the model not only learn new knowledge (i.e. plasticity) but also remember the knowledge from previous tasks.
@@ 6.0.1 Limitations and Future Works
Acknowledgement
This work is partially supported by the Army Research Office/Army Research Laboratory via grant W911-NF-20-1-0167 to Northeastern University, National Science Foundation CCF-1937500.